OpenStudy (anonymous):

I know this limit should be zero but i'm not sure why. \[\lim_{n \rightarrow \infty} \frac{ 917^(n+100) }{ 955^n }\]

OpenStudy (anonymous):

Just to clarify that is 917^(n+100) on the top

OpenStudy (anonymous):

Just going to throw a joke out here because it's 4:04am... "The limit does not exist!" - Mean Girls

OpenStudy (anonymous):

That's not really the answer though :P

OpenStudy (ash2326):

\[\frac{917^{n+100}}{955^{n}\] \[n-> \infty\] \[{\frac{917}{955}}^n \times 917^100\] \[n\to \infty\] so \[{\frac{917}{955} }^n->0\] \[0\times 917^100=0\]

OpenStudy (ash2326):

oops \[0\times 917^{100}=0\]

OpenStudy (anonymous):

Think i've got it, how i worked it out was like this. Could be completely wrong but oh well. \[\frac{ n(\frac{ 917^{n+100} }{ n }) }{ 955 }\] \[\frac{ 0 }{ 955 }= 0\] ?

OpenStudy (anonymous):

wait, left something out.

OpenStudy (anonymous):

ahhh, just imagine the bottom line as n(955)

OpenStudy (ash2326):

But this part \[\frac{917^{n+100}}{n}\ne 0\] It's of the form \[\frac{\infty}{\infty}\]